In a farming community 30% of the farmers grow oranges only, 10% of the farmers grow lemons only and 4% grow both lemons and oranges .what proportion of the farms in the community grow either lemons or oranges?

i.0.14
ii.0.24
iii.0.48
iv.0.34
v.0.44
1
Expert's answer
2021-09-02T07:25:23-0400
"P(O\cup L)=P(O-L)+P(L-O)+P(O\cap L)"

"=0.3+0.1+0.04=0.44"



v. 0.44